Discuss best arrangement of a set of ideal reactors (small mixed flow reactor, large mixed flow reactor and plug flow reactor) connected in series for any nth order reaction,n>0.


1
Expert's answer
2021-09-02T02:32:40-0400

The mixed reactor is always bigger than the plug flow reactor for any given job, and all positive reaction commands. The volume-to-volume ratio rises as response order rises. When the conversion is minimal, flow type has only a little impact on reactor performance. At high conversion, the performance ratio rises extremely quickly; as a result, a good depiction of the flow becomes critical in this range of conversion. The fluctuation in density throughout the reaction affects the design, although it is usually secondary to the variance inflow type.

Consider a Plug Flow Reactors in Series

N plug flow reactors

"X_1, X_2, . . . , X_N"

The fractional conversion of component A leaving reactor 1, 2, . . . , N

For the ith reactor

"\frac{V_i}{F_0}= \int_{X_{i-1}}^{X_i} \frac{dX}{-r}"

For several reactors with a total volume of V

"\frac{V}{F_o}= \sum^N_{i=1} \frac{V_i}{F_o}= \frac{V_1+V_2+...+V_0}{F_o}\\
=\int_{X_{0=0}}^{X_1} \frac{dX}{-r}+\int_{X_{1}}^{X_2} \frac{dX}{-r}+...+\int_{X_{N-1}}^{X_N} \frac{dX}{-r}=\int_{0}^{X_N} \frac{dX}{-r}"

N plug flow reactors in series with a total volume V gives the same conversion as

a single plug flow reactor of volume V